OpenStudy (anonymous):

Convert the numeral to a numeral in the base indicated. 16,629 to base 13

OpenStudy (valpey):

What are the first few powers of 13?

OpenStudy (ash2326):

decimal no.=16, 629 base 13,so the no. will have digits from 0 to 12

OpenStudy (anonymous):

you are french right? 13^x = 16.629 x = log(16.629)/log(13) x= 1.09599 \[13^{1.096} = 16.629\]

OpenStudy (anonymous):

i don' know

OpenStudy (anonymous):

trust me my answer is right

OpenStudy (anonymous):

went to MIT

OpenStudy (valpey):

Yes, but the key to converting to bases is a division and remainder problem. \[13^0 = 1;\ 13^1 = 13;\ 13^2 = 169;\ 13^3 = 2197;\ 13^4 = 28,561\]

OpenStudy (anonymous):

ok thanks

OpenStudy (valpey):

@zscdragon is trolling you.

OpenStudy (valpey):

16,629 is not greater than or equal to 13^4 so we won't be needing that one. How many times can 13^3 go into 16,629?

OpenStudy (ash2326):

we have to continuously divide the no. by 13 to get the number in base 13 1279 ______________________ 13|16629 16627 remainder =2 this is the first digit now divide 1279 by 13 we get 98 as quotient and 5 as remainder so it is the second digit Do you get till here?

OpenStudy (ash2326):

@85295james ??

OpenStudy (anonymous):

o

OpenStudy (ash2326):

huh?

OpenStudy (anonymous):

@85295james write my stuff it's the correct one ;)

OpenStudy (valpey):

Like I said, @zscdragon is a troll. @ash2326 's method is correct, even if it is confusing. The solution will be a four digit number in base 13. Let's call it ABCD where: \[A*13^3+B*13^2+C*13^1+D*13^0 = 16,629\]

OpenStudy (ash2326):

@zscdragon what's 8 in base 8?

OpenStudy (anonymous):

16629 / 13 = 1279 1279 * 13 = 16627 that means the las digit is 16629 - 16627 = 2

OpenStudy (anonymous):

16629 in base 13 is 7752

OpenStudy (anonymous):

relax people, he's french, i know what he meant it's not the same mathematical vocabulary ;)

OpenStudy (anonymous):

that was a log problem

OpenStudy (anonymous):

how do i List all the subsets of {1}.

OpenStudy (anonymous):

@85295james have you understood what i messaged you?

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

be clear.

OpenStudy (anonymous):

o yes i got it

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

but how do i List all the subsets of {1}.

OpenStudy (anonymous):

is it 7752..?